### v1

# NICE: To Optimize In-Context Examples or Not?

Abstract
Recent works have shown that large language models (LLMs) work remarkably well on a wide range of tasks through in-context learning and optimization of in-context examples (ICE). However, most of these studies assume either a fixed or no instruction provided in the prompt, leading to the apparent consensus that the optimization of in-context examples is critical for better performance. We challenge this consensus for instruction-tuned LLMs by investigating the necessity of optimizing in-context examples when task-specific instructions are provided, and find that there are tasks for which various ways of optimizing in-context examples yield diminishing returns. We introduce a task-specific metric called Normalized Invariability to Choice of Examples (NICE) that quantifies the learnability of tasks from a given instruction, and provides a heuristic that helps decide whether to optimize for instructions or ICE for any new task. On a wide range of tasks and a systematically created instruction set with gradually added details, we validate our hypothesis empirically by computing NICE with query-dependent bins of examples, comparing different instructions with ICE selection methods, and performing label perturbation experiments. We conclude that tasks can be divided into two broad classes based on the NICE metric, where the returns on ICE optimization follow predictable trends when instructions are provided in the prompt.
1 Introduction
The scaling up of large language models (LLMs) in terms of model size, training data, and compute in recent times has led researchers to notice their emergent in-context learning ability across a wide range of tasks (Brown et al., 2020; Radford et al., 2019; Kaplan et al., 2020). The in-context learning paradigm studies the ability to learn a new task based on an input context augmented with a few example demonstrations. The selection of good in-context examples (ICE) from a large pool of candidate training examples is a non-trivial optimization problem over a large search space (Liu et al., 2021).
The emergence of in-context learning ability in LLMs has motivated a growing body of literature on how to select the ICE to be added to an LLM’s input prompt (Liu et al., 2021; Lu et al., 2021; Zhang et al., 2022b; Qin et al., 2023). These papers propose various ICE selection techniques that can significantly improve LLMs’ performance across a range of tasks, leading to the apparent consensus in literature that optimizing ICE is critical for better in-context learning performance. Thus, it appears that improving the optimization of ICE is the key to better in-context learning using LLMs.
However, with the advent of instruction-tuning Wei et al. (2021); Chung et al. (2022); Ouyang et al. (2022), LLMs are also capable of following natural language instructions, making it possible to solve a task in a zero-shot manner, i.e., learning with just the instruction and no ICE. A natural question that follows is how instructions and ICE interact and whether they can be jointly optimized. Would optimization of ICE still matter after being given a descriptive or detailed task instruction? And more practically, given a task, what is better—optimizing the instruction or optimizing the ICE?
Using state-of-the-art LLMs, our main finding is that the returns on ICE optimization diminish as more detailed instructions are added to the input prompt. We find that a detailed instruction with randomly chosen ICE surpasses or matches the accuracy of a prompt with ICE selected using criteria from past work Liu et al. (2021); Levy et al. (2022); Rubin et al. (2021); Sorensen et al. (2022), across diverse classification tasks such as SST-2 and SST-5 for 2 and 5-way sentiment classification respectively Socher et al. (2013), MNLI for natural language inference Williams et al. (2017), TREC for 6-way question classification Li and Roth (2002), MedQA Jin et al. (2021) and FinQA for domain-specific question-answering Chen et al. (2021). Moreover, given a detailed instruction, even using ICE with incorrect labels provides the same accuracy as using the correct ICE on most tasks. On the other hand, there exists a different class of tasks, especially generative tasks such as MTOP Li et al. (2020) for semantic parsing and Break Wolfson et al. (2020) for Question Decomposed Meaning Representation (QDMR) generation, where ICE selection actually matters: a prompt containing the best chosen ICE outperforms a prompt containing a descriptive instruction with random ICE.
Therefore, contrary to prior work, ICE optimization per query may not always be the most effective strategy for increasing task performance, especially as task instructions are much simpler to write (e.g., using our proposed template). To characterize the tasks where ICE optimization may not matter, we introduce a metric that measures the invariability of an LLM’s task performance to ICE from different bins based on a query-dependent scoring function (e.g., cosine similarity). The metric, called Normalized Invariability to Choice of Examples (NICE), partitions candidate training examples into bins based on the above scoring function, and compares the performance of the best bin for picking random ICE with the mean performance across all bins. We show that this simple metric can effectively capture the dependence of tasks on ICE optimization. Tasks such as MTOP and BREAK score low on NICE metric (<0.4) whereas all the classification tasks score high (>0.85). Further, consistent with our results above, the NICE score increases as more detailed instructions are introduced in the prompt.
Figure 1: An overview of prompt optimization by improving the instruction and ICE following the NICE metric based for any new task following our hypothesis.
As a practical implication, our work provides a method to decide the best option for improving an LLM’s task accuracy: ICE optimization or improving the instruction (see Figure 1). Given a task, we provide a template for creating a detailed instruction. Then, given the task and an instruction from the template, we provide an efficient procedure to compute the NICE score. If the NICE score is high, then it may be more effective to try improving the instruction; ICE optimization may not add much value. However, for tasks where it is low, ICE optimization will be effective.
To summarize, we make the following key contributions:
  • Contrary to prior work, we show that ICE optimization does not always lead to performance improvement for LLMs. Across a range of tasks, we find that the performance of prompts improves with better-quality instructions, and this improvement remains constant even when true ICE are replaced with ICE with incorrect labels.
  • For practical usage, we introduce a task-specific metric called NICE that can predict which tasks will benefit from ICE optimization.
  • For tasks with high NICE score, we find that detailed instructions with random ICE obtain comparable or better performance than prompts with optimized ICE. Conversely, for tasks with low NICE score, prompts with optimized ICE obtain better performance.
In this study, we consider a prompt to contain both the instruction and the in-context examples, and compare the effectiveness of these components on a number of tasks. Previous works Min et al. (2022a); Wei et al. (2023) have argued that LLMs benefit from ICE even though they may not actually learn input-label mappings from ground truth examples. We find that the benefit of ICE changes with detailed instruction. There is a large class of tasks for which ICE optimization gives diminishing returns. Specifically, with the addition of a reasonable instruction in the prompt, dependence on ICE reduces significantly.
Dataset Type of Task Train Size Test Size NICE (NI → DI) Evaluation
SST-2 2-class Sentiment CF 10K 1K 0.99 → 1.00 Accuracy
MNLI Textual Entailment 10K 1K 0.97 → 0.99 Accuracy
MedQA Multi-choice QA 10K 1K 0.98 → 0.98 Accuracy
TREC 6-class Question CF 5.9K 0.5K 0.52 → 0.98 Accuracy
FinQA Hybrid QA 10K 1K 0.95 → 0.96 Accuracy
SST-5 5-class Sentiment CF 8.53K 1K 0.91 → 0.94 Accuracy
NL2BASH Code Generation 10K 0.6K 0.85 → 0.85 BLEU Score
MTOP Semantic Parsing 10K 1K 0.29 → 0.38 Exact Match
Break Question Decomposition 10K 1K 0.43 → 0.20 Graph Edit Dist.
Table 1: Details of the task datasets used in our evaluation. Tasks sorted on NICE with instructions with NL2BASH, MTOP, and Break being the low-NICE tasks. Abbreviation key: NI (No Instruction), DI (Detailed Instruction), CF (Classification). Note that the metric is not centered around $0.5$, and a NICE score of $0.85$ indicates a significant scope of improvement with better examples.
2 Related Works
Several related works have studied the optimization of both instructions and ICE for LLMs on various classification and generation tasks.
2.1 In-context Example Selection Methods
Learning-free selection methods: Liu et al. (2021) propose a kNN-based method for selecting the top-$k$ candidates ranked in the reverse order of their cosine-similarity with the query based on the assumption that in-context examples with high similarity in the embedding space would help the most on downstream tasks. Sorensen et al. (2022) use a mutual-information-based metric for comparing two prompts and select the one that maximizes mutual information with the query. Gonen et al. (2022) attempt to select in-context examples with low perplexity, while Levy et al. (2022) show that diverse demonstrations improve the compositional generalization ability of LLMs.
Learning-based selection methods: Rubin et al. (2021) propose a two-stage retrieval method called Efficient Prompt Retrieval (EPR) to select ICE. For a specific input instance, they utilize a BM-25-based retriever to select candidate examples and subsequently employ a retriever trained through contrastive learning. Li et al. (2023) further enhance EPR by proposing a Unified Demonstration Retriever (UDR) across different tasks. Lu et al. (2022) propose a policy-gradient method for selecting in-context examples using a specific policy formulation for a given query. Ye et al. (2023) retrieve the complete set of demonstrations rather than extracting individual ones in order to capture inter-demonstration relationships. They used contrastive learning to train a Determinantal Point Process (DPP)-based retriever to match the scores of the language model output, obtaining the most likely demonstration set through maximum a-posteriori during inference.
2.2 Jointly Optimizing In-context Examples and Instructions
Some works (Zhou et al., 2022; Pryzant et al., 2023) use an LLM to improve the instruction alone but do not consider examples. Zhang et al. (2022a) introduce TEMPERA, an RL-based optimization scheme using discrete instruction and example sets. However, the search space for unconstrained joint optimization becomes exponentially large for large instruction and example spaces. While these works showcase the importance of optimizing both instructions and examples for various tasks, they do not perform a comprehensive comparison of the two. Our objective instead is to compare the two methods of task-learning for various tasks and evaluate how learning from instructions versus examples varies across tasks. We empirically show that the choice of optimal task-learning paradigm depends on the task, and we suggest a method to analyze any given task to gain some traction on which of these optimizations is expected to benefit model performance.
3 Formulating the Task-specific Metric
We propose a task-specific metric designed to categorize language tasks into two types based on their learnability from instructions and in-context examples. First, we employ a grouping function $f$ to partition the training set of candidate examples into multiple bins based on the query. Subsequently, for each bin, we compute the model’s performance by selecting random ICE from that bin for a set of sampled queries to get a score for each bin denoted by:
$S(j,\text{I},\mathcal{D})=\mathop{\mathbb{E}}\limits_{x\sim\mathcal{D}}\mathop{\mathbb{E}}\limits_{e_{j}\sim b^{x}_{j}}\left[Y_{LM}(x|(\text{I}+e_{j}))\right],$ (1)
where $Y_{LM}$ is the performance measure (higher is better) of the language model, $\mathcal{D}$ is the set of test queries, I is the given instruction and $b^{x}_{j}\in\mathcal{B}$ is the $j$-th bin for query $x$. Next, we look at a few intuitive requirements from such a metric and finally define the NICE metric in Subsection 3.2.
3.1 Metric Requirements
Here we outline some intuitive requirements and considerations while formulating a metric that captures the non-benefit of ICE optimization for a given task and a given instruction:
  1. 1.
    Property 1: The metric should be bounded within (0,1], establishing a well-defined baseline to compare all tasks.
  2. 2.
    Property 2: It should be linearly invariant to the performance measures for each task, ensuring independence from variations in model capability and task difficulty.
  3. 3.
    Property 3: If $S(j,\text{I})\approx\max\limits_{j}S(j,\text{I})$ for all $j$, i.e., the performance of the model is invariant to the bin from which the in-context examples are chosen, the metric should be close to $1$.
  4. 4.
    Property 4: Let $b_{\text{max}}=\arg\max_{j}S(j,\text{I})$ be the best bin. If, for all $j\neq b_{\text{max}}$, $S(j,\text{I})\ll S(b_{\text{max}},\text{I})$, this represents a strictly retrieval task where the budget needs to be spent on in-context example optimization. In this scenario, the metric should approach 00.
Since API calls to LLMs are both costly and time consuming, another requirement is that the metric should be computable much more efficiently than typical ways to optimize over the search space of all candidate examples.
To compare different instructions, we can define a partial order on the set of instructions as $I\preceq I^{\prime}$ (in words, $I$ is Pareto-dominated by another instruction $I^{\prime}$) if $S(j,I)\leq S(j,I^{\prime})$, $\forall j$. One can prove (under simple assumptions) that if there exists an $I^{*}$ that is not Pareto-dominated by any other $I^{\prime}$, then it must be a maximizer of $\sum_{j}w_{j}S(j,I)$ for some non-negative weights $w_{j}$’s (Harsanyi, 1955). The weights $w_{j}$’s are typically uniform or chosen by domain experts (Harsanyi, 1955; Gale, 1960). For simplicity, we use uniform weights to define our metric. In the next section, we introduce the NICE metric, and in Appendix A.2, we prove that it follows all these requirements, and discuss additional benefits of its non-uniform variants.
3.2 Normalized Invariability to Choice of Examples (NICE)
Our metric, called the Normalized Invariability to Choice of Examples (NICE) of a task for a given instruction, is defined as the ratio of the expected score over all bins to the maximum amongst them. Mathematically, NICE is defined as follows:
$\text{NICE}(\mathcal{D,\text{I}})=\frac{\mathop{\mathbb{E}}\limits_{0<j\leq|\mathcal{B}|}\left[S(j,\text{I},\mathcal{D})\right]}{\max\limits_{0<j\leq|\mathcal{B}|}S(j,\text{I},\mathcal{D})},$ (2)
where $S(j,\text{I})$ is the score for the $j$-th bin in $\mathcal{B}$ for a given instruction. Note that this score can be computed for any choice of task, any instruction, any performance measure $Y_{LM}$, and any grouping function $f$ for partitioning the candidate examples. Liu et al. (2021) show that cosine similarity with the query is a good proxy for selecting in-context examples. Thus, we set up our baseline to study the performance trend in various datasets by binning the examples ($\mathbf{x}$) based on their similarity score with the query ($\mathbf{q}$). We assume that the grouping function $f$ to be as follows:
$f:=sim(\mathbf{x},\mathbf{q})=\frac{\mathbf{x}^{\top}\mathbf{q}}{\lVert\mathbf{x}\rVert\lVert\mathbf{q}\rVert}.$ (3)
We partition the candidate examples into bins based on the value of the grouping function $f$ for example $\mathbf{x}$, e.g., $0-10\%$, $10-20\%$. Combining Eqs. (1) and (2) with the grouping function from Eq. (3) gives us the NICE baseline we set up to distinguish tasks.
We hypothesize that for tasks with a high NICE, the task is learnable from the instructions and does not require carefully selected examples. On the other hand, for tasks with a lower NICE, examples play a major role in task learning, so we suggest spending more budget on optimizing ICE. A simpler task (such as SST-2) can be expected to have a flat trend in performance across bins, and thus a high NICE. On the other hand, for complex tasks where the model needs to understand the input-output relationship and its syntax (such as MTOP, Break and NL2BASH in our study) by looking at the retrieved in-context examples, we expect to have a lower NICE. Our simple baseline using cosine similarity-based bins can be easily extended to consider more advanced clustering strategies by using heuristics such as influence functions (Nguyen and Wong, 2023).
We experiment with the following datasets in line with other works in in-context learning: Stanford Sentiment Treebank (SST-2 and SST-5) (Socher et al., 2013), question multi-class classification in TREC (Li and Roth, 2002), textual entailment in MNLI (Williams et al., 2017), semantic parsing in MTOP which is a multilingual task-oriented semantic parsing dataset with hierarchical intent-slot schema.(Li et al., 2020), question decomposition in Break where input is a complex question (x) and the task is to map the question to its QDMR steps ($\mathbf{s}=\langle s_{1},...,s_{n}\rangle$) (Wolfson et al., 2020), code generation in NL2BASH, and hybrid and multi-choice question-answering in FinQA (Chen et al., 2021) and MedQA (Jin et al., 2021). We download standard versions of all these datasets from HuggingFace (Delangue and team., 2016). Details about our tasks along with their NICE with and without instructions are given in Table 1.
Figure 2: Per-bin performance trend of various datasets with distance-based bin using GPT-4. Arrow indicates change in NICE when shifting from no instruction to a detailed instruction. Note that the y-axis shows the ratio of performance of a given bin with the max performance over all bins with no instructions.
Figure 3: Comparison of in-context example optimization methods against instructions for various classification and structured generation tasks using GPT-3.5-turbo. Abbreviation key: LM (Task Def. with Label Meanings), R (Task Def. with Rules). All ICL methods pick $8$-shot examples.
4 Experiment Design
4.1 Prompt Structure and Instruction Set
For a fair comparison across instructions and easy applicability to any new task, we use a standard prompt template that consists of an instruction, $4$-shot demonstrations $(x_{i},y_{i})$ from a training set $\mathcal{D_{T}}$, and the query $x$ from the test set with a cue for the model to predict the corresponding $y$. We define a consistent structure to create a general set of instructions for all tasks, iteratively adding useful, task-specific information to the instruction starting with no instruction following a guideline similar as in Zhang et al. (2023). A consistent template that does not involve human creativity ensures that our findings are not constrained to a narrow distribution of carefully-crafted prompts. We use this gradation of adding detailed information as a rough proxy for "improving" the instruction based on results from Zhang et al. (2023).
For classification tasks, we consider the following kinds of instructions:
  • No Instruction (NI): We provide no instruction and the model is expected to learn only from in-context examples.
  • Task Definition (TD): A simple task definition without any other information.
  • Task Definition with Label Space (+LS): Including the label space in the task definition. (e.g. [“positive”, “negative”] for a binary sentiment classification task.
  • Task Definition with Label Meanings (+LM): We specify the meanings of the labels in the label space along with the task definition.
  • Delusive Instruction (DI): We also evaluate the model’s task learnability when the instruction is misleading, i.e., describing a task with shuffled labels. (e.g. “positive” for “negative”).
For generative tasks, we experiment with the ‘No-Instruction’ and ‘Task-Definition’ cases in the same way as defined for classification tasks. Since there is no notion of label space in generation tasks, we introduce an additional variety of detailed instruction:
  • Task Definition with Rules: A simple task definition accompanied with syntactic rules for the task.
Figure 4: Label choice ablation (gold to random) performance for classification and structured generation tasks using GPT-3.5-turbo. Abbreviation key = DI (Delusive Instruction), NI (No Instruction), TD (Task Definition), +LS (Task Def. with Label Space), +LM (Task Def. with Label Meanings), +R (Task Def. with Rules). Note that classification tasks do not have rules while generation tasks do not have label spaces and delusive instructions.
4.2 Investigations
We structure our study around the following research questions:
RQ1: What kinds of tasks require in-context example optimization?
To investigate whether ICE optimization is needed for a given task, we evaluate various tasks based on the NICE metric defined in Eq. (2).
Setup: We assume that the training split of each dataset acts as a pool of candidate examples, and pick 50 queries randomly from the test split and partition the training set for each query into groups/bins based on the grouping function $f$ defined in Eq. (1). We randomly sample 10 sets of 4-shot demonstrations from a specific bin for a given query and evaluate the model’s performance and report the averaged out results over 50 queries for each group. For better visualization, we scale the accuracies across the different groups by dividing them by their mean. This scaling of accuracy values doesn’t affect the trend as our metric is invariant to linear scaling (see Subsection 3.1).
RQ2: How do different kinds of in-context example selection approaches fare against a detailed instruction?
We consider three different baselines for methods to select in-context examples for a give query based on different kinds of properties:
Top-$k$-BERT: This dense retrieval method uses BERT (Devlin et al., 2018) embeddings to compute the cosine similarity of examples with the query and selects the top-$k$ examples.
BM-25: BM-25 is a sparse bag-of-words retrieval algorithm that ranks examples using a tf-idf -based relevance score.
DPP: Determinantal Point Processes (DPP) Borodin (2009) computes a probability measure over the possible subsets of a discrete set. They model the repulsion between the items in a chosen subset by considering the volume of the parallelepiped formed by the feature vectors of the items in the chosen subset as a measure of probability of picking up that subset, as of consequence of which, the likelihood of selecting a subset decreases as items become more similar. Therefore, DPP ensures that the chosen subset is diverse in terms of specified criteria, such as representativeness.
Using each of the above methods, we pick up the top $k=8$ examples as the in-context examples with the instruction, similar to other works that use $k\in\{4,8,16\}$ (Min et al., 2022a; Wei et al., 2023; Lu et al., 2021; Zhang et al., 2022b).
RQ3: How much do the ground-truth labels of ICE matter as we keep on improving the instruction based on a fixed template?
Min et al. (2022b) identify four aspects of demonstrations which serve as a learning signal to the LLM: the input-label mapping, the distribution of the input text, the label space and the format. Concretely, they say that in-context examples play a role by informing the model of the label space or the format of the output. Pan et al. (2023) discuss the phenomena of task learning and task recognition as the two ways through which an LLM learns from demonstrations by comparing the performance of randomly labeled examples against abstract and gold-labeled ones.
Relatedly, in this work, we evaluate the role of input-label mapping for in-context learning. Specifically, we investigate whether, instead of relying on the demonstrations for information about the label space or format of the output, it is possible to subsume this information in the instruction such that this leads to a minimal drop in performance. To this extent, we introduce perturbations in the in-context examples by randomizing the ground-truth labels of the in-context examples. In the case of classification tasks, we assign randomly chosen labels $y\sim\mathcal{Y_{T}}$ to the inputs of the in-context examples, where $\mathcal{Y_{T}}$ denotes the label space of the output for a particular task $\mathcal{T}$. In the case of generation tasks, we shuffle the ground truth values $y$ of the in-context examples amongst themselves.
5 Results
We devise three experiments to answer the three research questions highlighted in 4.2 and present the results for them in this section. We experiment with state-of-the-art LLMs with a good instruction following ability, and use GPT-4-Turbo for computing the NICE score for each task, and GPT-3.5-Turbo for all other experiments. We also show NICE results from open-source models in Appendix A.3.
5.1 High-NICE and Low-NICE Tasks (RQ1)
Fig. 2 shows the trend in performance with and without instructions using GPT-4-Turbo:
  1. 1.
    No Instruction: When we prompt the model with just in-context examples without any instruction, we observe the NICE for relatively simpler tasks (SST2, SST5, and MNLI) to be higher than more complex tasks (TREC, Break, NL2BASH, and MTOP).
  2. 2.
    Detailed Instruction: When a detailed task-specific instruction is added to the prompt along with in-context examples, we observe a flattening effect of the instructions for TREC, SST5, and MTOP, thus increasing the NICE for these tasks. This evidence supports our hypothesis that the effect of in-context example optimization decreases in the presence of instructions.
We observe in Fig. 2 that for tasks with a high NICE, it is not helpful to selectively retrieve in-context examples conditioned on the query.
As described in 2, Rubin et al. (2021) propose a contrastive-learning objective for selecting in-context examples. Specifically, they define a score function for each of the examples in the set of candidates given by:
$s_{ij}=P_{LM}(y_{i}|e_{ij},x_{i})\quad\forall j\in\{1,2,...,L\}\quad\forall i$
In their case, the contrastive learning objective plays the role to increase the relevance of the subset of examples with higher LM scores over the others for a given query. We show that in a scenario, where we have example equivalence in terms of model’s performance across the different candidate examples, even this exercise doesn’t lead to an improvement in performance.
5.2 Optimizing In-context Examples: A Comparison with Instructions (RQ2)
Fig. 3 shows a comparison between the performance of different in-context example selection methods with no instructions against randomly sampled examples with a detailed task-specific instructions. We find that for high NICE tasks (SST2, SST5, TREC, MNLI, FinQA, and MedQA) which possess high-instruction learnability, the performance with even randomly sampled in-context examples complemented with a carefully crafted task-specific instruction surpasses or matches the performance of the above mentioned in-context example selection methods with no instruction.
On the other hand, for tasks with a low NICE (MTOP, Break, and NL2BASH), we see that in-context example optimization helps in improving the performance over randomly sampled examples accompanied with instructions. Interestingly, we observe that the the distance of the task’s NICE with complete instruction learnability (i.e., NICE =1) is proportional to the difference in performance of in-context example optimization and instructions. For instance, for tasks such as NL2BASH and SST5, we see a slight dip in performance when using instruction, while the dip is significant for MTOP and Break.
5.3 Label Perturbation (RQ3)
In Fig. 4, we see that as we move to more detailed instructions, the difference in performance with gold-labeled examples against randomly-labeled examples keeps on diminishing and ultimately becomes negligible for detailed task-specific instructions. This validates our hypothesis that for tasks with a high NICE, good instructions suffice and the dependence of the performance on examples is less and it reduces further with better instructions. (see 4.1 for the instruction templates). Thus, we see that a better instruction helps in reducing the model’s reliance on the in-context examples to understand input-label mapping. In the case of classification tasks, there is a drop in performance with the incorporation of a delusive instruction in the prompt; this drop is also an indicator of the model’s dependence on the instruction to understand a task.
Finally, on tasks with a low NICE (MTOP, Break and NL2BASH), we see a different trend in Fig. 4. While with no instructions we see a significant improvement in the performance ($\sim$1.04x for MTOP and $\sim$1.25x for Break), the improvement from gold-labeled examples stays consistent with the addition of task definitions and the syntax rules of the task. Furthermore, the narrowing of the performance gap between gold and random labels for in-context examples can also be observed in the case of these tasks, albeit at a slower rate compared to tasks with a high NICE. Consequently, the importance of optimizing in-context examples becomes more evident in this scenario.
6 Conclusion
In this work, we challenge the assumption that optimizing in-context examples is always crucial for better performance of LLMs on downstream tasks. Through extensive experimentation on instruction-tuned models, we demonstrate empirically that ICE optimization yields diminishing returns in the presence of task-specific instructions for a class of tasks. We introduce Normalized Invariability to Choice of Examples (NICE) as a metric to analyze the learnability of tasks from given instructions, and show that for hig-NICE tasks, better-quality instructions lead to comparable or better performance than ICE optimization baselines. We also conduct label choice ablation experiments, confirming that with detailed instructions, the model’s performance becomes less sensitive to the choice of labels in the in-context examples. We hope that our work sheds some light on the effects of instructions and examples for prompting LLMs and would lead to a more systematic method to optimize prompts for in-context learning of any new task.
7 Ethics Statement
We aim to advance the field of machine learning by improving our understanding of how universal large language models (LLMs) learn various tasks from instructions and prompts. We use proprietary models (GPT-4-Turbo, GPT-3.5-Turbo) accessed through OpenAI’s API endpoints and publicly available datasets from HuggingFace. Our work can help the scientific community evaluate new tasks and guide them toward building better prompts, which can have performance impacts while testing and deploying LLMs for downstream tasks. Like any other work advancing the field of machine learning, our work could also help understand and improve prompts for negative tasks, but we strongly discourage any negative usage extending from our work and LLMs.
Appendix A Appendix
A.1 Task-wise Instruction Sets
A.1.1 Classification Tasks
In 4.1, we discuss the structured instruction sets we use. In this section, we give some concrete examples. We also note that all of these instructions follow standard templates and none of them required human creativity. The total amount of time it took to create all of them for $9$ datasets was under $1$ hour. We show the instruction sets for SST-2 and MTOP, and refer to our codebase for all other datasets.
Figs. 6, 7, 8, and 5 show the task definition (TD), TD with label space, TD with label meaning, and delusive instruction for SST-2.
{promptbox}
Your task is to classify the given movie review based on the sentiment expressed. If the overall opinion of the reviewer based on the keywords or phrases in the review is positive, classify it as "negative". If the overall opinion of the reviewer based on the keywords or phrases in the review is negative, classify it as "positive". Below are some demonstrations of the task, look at the examples to understand the task.
Figure 5: Classification Task - Delusive Instruction.
{promptbox}
Your task is to classify the given movie review based on the sentiment expressed. Below are some demonstrations of the task, look at the examples to understand the task.
Figure 6: Classification Task - Task Definition.
{promptbox}
Your task is to classify the given movie review based on the sentiment expressed, into "positive" or "negative". Below are some demonstrations of the task, look at the examples to understand the task.
Figure 7: Classification Task - Task Definition with Label Space.
{promptbox}
Your task is to classify the given movie review based on the sentiment expressed. If the overall opinion of the reviewer based on the keywords or phrases in the review is positive, classify it as "positive". If the overall opinion of the reviewer based on the keywords or phrases in the review is negative, classify it as "negative". Below are some demonstrations of the task, look at the examples to understand the task.
Figure 8: Classification Task - Task Definition with Label Meaning.
A.1.2 Generation Tasks
Fig. 9 and 10, we illustrate the task definition (TD) and TD with rules for MTOP:
{promptbox}
You will be given a user utterance in a specific domain and a particular language. Your task is to convert that utterance into a logical form representation. Below are some demonstrations of the task, look at the examples to understand the task and answer the query at the end.
Figure 9: Generation Task - Task Definition.
You will be given a user utterance in a specific domain and a particular language. Your task is to convert that utterance into a logical form representation. To do so, you need to abide by the following rules:
1. Identify the intent of the user from the utterance depending upon the domain.
2. Tag the user utterance with slot labels depending upon the domain. Slot label values can be text spans from user-utterance or nested queries.
3. The standard format of the output is:- [IN: <user-intent> [SL: <slot-label-1> <slot-label-value-1> ] [SL: <slot-label-2> <slot-label-value-2> ]…[SL: <slot-label-n> <slot-label-value-n> ] ] if there are n slot labels tagged in an utterance.
4. In some cases, the slots can be nested with intents within them, for those cases, use this format:- [IN: <user-intent> [SL: <slot-label> [IN: <user-intent> [SL: <slot-label-1> <slot-label-value-1> ] [SL: <slot-label-2> <slot-label-value-2> ]…[SL: <slot-label-n> <slot-label-value-n>]]]]
5. Apply the same grammar rules in case of nested queries for the inner level.
Below are some demonstrations of the task, look at the examples to understand the task and answer the query at the end.
Figure 10: Generation Task - Task Definition with Rules.
The “No Instruction” case for both classification and generation tasks has an empty instruction appended by the examples. For instruction templates for all our classification and generation tasks, we refer the interested reader to our codebase.
A.2 Proposition Proofs and Discussion on the NICE Metric
In this section we prove that the NICE metric defined in Section 3 indeed follows the propositions in Section 3.1, and discuss some ways the NICE score can be used for prompt optimization.
Proposition 1: Boundedness of NICE
The NICE metric is bounded within (0,1], establishing a well-defined baseline to compare all tasks.
Proof: Since NICE is defined as the ratio of the expected score over all bins to the maximum amongst them, it can be expressed as:
$\text{NICE}(\mathcal{D,\text{I}})=\frac{\mathop{\mathbb{E}}\limits_{0<j\leq|\mathcal{B}|}\left[S(j,\text{I},\mathcal{D})\right]}{\max\limits_{0<j\leq|\mathcal{B}|}S(j,\text{I},\mathcal{D})}$
The expectation of a bounded variable is itself bounded in (0,1], thus establishing the boundedness of NICE.
Proposition 2: Invariance to Linear Scaling
NICE should be linearly invariant to the performance measure, ensuring independence from linear variations in model capability and task difficulty.
Proof: Consider the definition of NICE where performance (and thus the scores $S$) scale linearly with a factor $k$ by linearity of expectation:
$\text{NICE}(\mathcal{D,\text{I}})=\frac{\mathop{\mathbb{E}}\limits_{0<j\leq|\mathcal{B}|}k.S(j,\text{I},\mathcal{D})}{\max\limits_{0<j\leq|\mathcal{B}|}k.S(j,\text{I},\mathcal{D})}$
The linear invariance comes from the fact that scaling the performance measure $Y_{LM}(x|(\text{I}+e_{i}))$ by a constant factor will not affect the ratio, as both the numerator and denominator will be scaled equally. Thus, NICE remains invariant to linear scaling.
Proposition 3: Invariance to Bin Selection
If $S(j,\text{I},\mathcal{D})\approx\max\limits_{j}S(j,\text{I},\mathcal{D})$ for all $i$, i.e., performance of the model is invariant to the bin from which the in-context examples are chosen, $\text{NICE}\to 1$.
Proof: Since, there is an equal probability to sample examples from every bin,
$\mathop{\mathbb{E}}\limits_{0<j\leq|\mathcal{B}|}\left[S(j,\text{I},\mathcal{D})\right]=\frac{\sum_{j=1}^{|\mathcal{B}|}S(j,\text{I},\mathcal{D})}{|\mathcal{B}|}$
If the model’s performance is agnostic to the bin from which the ICE are picked, formally
$\forall jS(j,\text{I},\mathcal{D})\approx\max\limits_{i}S(j,\text{I},\mathcal{D})$
$\frac{S(j,\text{I},\mathcal{D})}{\max\limits_{j}S(j,\text{I},\mathcal{D})}\approx 1$
$\text{NICE}=\frac{\mathop{\mathbb{E}}\limits_{0<j\leq|\mathcal{B}|}S(i,\text{I},\mathcal{D})}{\max\limits_{0<j\leq|\mathcal{B}|}S(j,\text{I},\mathcal{D})}\approx\frac{|\mathcal{B}|}{|\mathcal{B}|}=1$
Proposition 4: Strict Retrieval Task Scenario
Let $b_{\text{max}}=\arg\max_{j}S(j,\text{I},\mathcal{D})$ be the best bin. If, for all $j\neq b_{\text{max}}$, $S(j,\text{I},\mathcal{D})\ll S(b_{\text{max}},\text{I},\mathcal{D})$, representing a strictly retrieval task where the budget needs to be spent on in-context example optimization. In this scenario: $\mathrm{\text{NICE}}\to 1/\lvert\mathcal{B}\rvert$ and $\lim\limits_{{\lvert\mathcal{B}\rvert\to\infty}}\mathrm{\text{NICE}}=0$.
Proof: In this case, let the maximum score be denoted by,
$m=S(b_{\text{max}},\text{I},\mathcal{D})$
If for $j\neq b_{\text{max}}$, $\frac{S(j,\text{I},\mathcal{D})}{m}\ll 1$. Assuming that we have a finite number of bins, the metric in this case reduces to
$\text{NICE}=\frac{\mathop{\mathbb{E}}\limits_{0<j\leq|\mathcal{B}|}S(i,\text{I},\mathcal{D})}{\max\limits_{0<j\leq|\mathcal{B}|}S(j,\text{I},\mathcal{D})}\approx\frac{m}{m.\lvert\mathcal{B}\rvert}=\frac{1}{\lvert\mathcal{B}\rvert}$
Now, as the number of bins approaches infinity (for instance, in a scenario where examples are segregated based on their absolute distance from the query rather than percentile), the granularity of the partitioning becomes infinitely fine. Therefore,
$\lim\limits_{{\lvert\mathcal{B}\rvert\to\infty}}\mathrm{\text{NICE}}=0$
The bin-wise NICE metric can also be used for instruction/prompt optimization. One possible way is as follows: Start with an initial instruction that maximizes $\sum_{j}S(j,I)$. Then, using weights $w_{j}\propto S(j,I)$, optimize $\sum_{j}w_{j}S(j,I)$ to find the next instruction, and iterate. This is almost an alternating optimzation over instruction and examples (if not examples, the weights over different bins from which examples are chosen) that intuitively optimizes instructions for the examples (or bins) that matter more for the task at hand. We leave this exploration and other uses of NICE for future work.
A.3 NICE Results for Open-Source Models
NICE of various datasets with and without instructions for “Mixtral” (an 8x7B mixture-of-experts model) are shown in Fig. and for “Llama-2-70b-chat-hf” are shown in Fig. 12. We note similar observations as GPT-4 when using open-source language models.
Figure 11: Per-bin performance trend of various datasets with distance-based bin using “Mixtral”, a mixture-of-experts model. Arrow indicates change in NICE when shifting from no instruction to a detailed instruction. Note that the y-axis shows the ratio of performance of a given bin with the max performance over all bins with no instructions.
Figure 12: Per-bin performance trend of various datasets with distance-based bin using “Llama-2-70b-chat-hf”. Arrow indicates change in NICE when shifting from no instruction to a detailed instruction. Note that the y-axis shows the ratio of performance of a given bin with the max performance over all bins with no instructions.
A.4 Pre-trained Models: Is Instruction-tuning the key?
While the results demonstrated in Figs. 2, 12, and 11 validate our hypotheses for instruction-tuned models, we see in Fig. 13 that they are valid for pre-trained models as well. Pre-training on next-token prediction enables models to learn a probability distribution $p_{LM}(\mathbf{x})$ over variable length sequences, thus allowing them to learn task-specific priors empowering them with zero-shot learnability, as was observed by Choi et al. (2022). We see that even models not explicitly fine-tuned on instruction-following exhibit the ability to learn from contexts.
We hypothesize that incorporating task-specific information containing label space information and other contextual cues provide substantial improvement for LLMs to predict the next tokens more accurately.
In Fig. 13, we show the results of the Llama-2-13B model on TREC (classification) and Break (generation) tasks. We see a partial pareto-dominance in case of semantic classification task like TREC (with high NICE) while it is not observed for the schema-based tasks like Break which has a low-NICE score.
Figure 13: Comparing different instructions on TREC and Break using pre-trained Llama-2-13B. Abbreviation key same as in Fig. 4.